Need to know how to work out this problem! 3x+y=8
-2x-y=6

Found 2 solutions by mananth, Alan3354:
Answer by mananth(16946)

3x+y = 8 .............1
-2x-y = 6 .............2
Eliminate y

Add the two equations
1 x = 14
/ 1
x = 14
plug value of x in (1)
3 x + 1 y = 8
42 + 1 y = 8
1 y = 8 -42
1 y = -34
y = -34

Answer by Alan3354(69443)

3x+y=8
-2x-y=6
-------- Add
x = 14
---------
Sub for x in the 1st eqn
3*14 + y = 8
y = -34